OpenStudy (swissgirl):

Show that the function of \( h: \mathbb{Z} \to \mathbb{Z} \) defined by h=3x is not a ring homomorphism

OpenStudy (swissgirl):

h(x+y)=3(x+y)=3x+3y=h(x)+h(y) So its homomorphic under addition

OpenStudy (swissgirl):

\(h(x*y)=3(x*y) \neq 3x*3y = h(x)*h(y) \)

OpenStudy (anonymous):

tada

OpenStudy (swissgirl):

YAYYYYYYYYYYY

OpenStudy (swissgirl):

SO that is all i gotta show?

OpenStudy (swissgirl):

I guessssssssssssss

OpenStudy (anonymous):

yesh